Session: Nov 2000 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Thu Nov 16 12:47:11 MET 2000) Minor failures
On two occasions a few scans lost due to unexpected stop of the tape on its own in the middle of a scan (at rougly the same position during reverse motion of tape). The lost observations are for 23:34 to 23:57 and 1:02 to 1:23 UTC. No other known problems.

Effelsberg (A. Kraus at Thu Nov 16 15:30:25 MET 2000) Minor failures
Scans between UT0613-0654 lost: problems with the VCs frequencies (human error).

Onsala (A. Polatidis at Fri Nov 17 12:57:05 MET 2000) Success

Jodrell_Bank (Alastair Gunn at Mon Nov 20 15:12:11 MET 2000) Minor failures
Lovell telescope late on first source due to MERLIN drive problems. On source at 23:06 UT. No other known problems.

Cambridge (Alastair Gunn at Mon Nov 20 15:12:29 MET 2000) Success
No known problems.

Noto (C.Trigilio at Mon Nov 20 16:37:25 MET 2000) Success

Medicina (C. Bortolotti at Tue Nov 21 12:02:08 MET 2000) Success

Westerbork (Tony Foley at Wed Nov 22 09:00:08 MET 2000) Minor failures
Some EMI. Otherwise no known problems